OpenStudy (anonymous):

Describe two ways to estimate the product 5 1/8 x 7 9/10 . Find the estimate for this product using each method.

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

5 1/8 = 41/8 and 7 9/10 = 79/10 So, calculate 41/8 x 79/10

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I am on a big timed test i just need the answer!

OpenStudy (anonymous):

The two methods I think are 1) First mutiply 41x79 and then divide it by 8x10. you will get a decimal as an answer. 2) first divide 41 by 8 and 79/10. you will get two decimals and add them.

OpenStudy (anonymous):

The answer you get is 40.4875 from both the methods